OpenStudy (anonymous):

Combine and simplify the following radical expressions sqrt 7-5/sqrt7+sqrt8=

OpenStudy (jdoe0001):

\(\bf \sqrt{7}-5\sqrt{7}+\sqrt{8}\) same radical terms, combine their coefficient